OpenStudy (anonymous):

Please help! A kicker kicks a football toward the opponent's goal line during a game. The ball begins its flight with an initial velocity of 100 feet per second when it is kicked at a height of 3 feet by the kicker. To the nearest foot, what maximum height will the ball reach? You may use the function: h(t)=-16t^2+100t+3 100 feet 129 feet 159 feet The answer cannot be determined

OpenStudy (anonymous):

Do you know how to use the quadratic formula?

OpenStudy (anonymous):

I think that's what you use

Mehek (mehek14):

\(\large{(\dfrac{-b}{2a},f(\dfrac{-b}{2a}))}\)

OpenStudy (anonymous):

i don't get it, what would i do after finding -b/2a?

Mehek (mehek14):

first find a and b

OpenStudy (anonymous):

a= -16 b=100 right?

Mehek (mehek14):

\(f(x)=ax^2+bx+c\)

Mehek (mehek14):

yes so plug it in

OpenStudy (anonymous):

25/8

Mehek (mehek14):

\(\dfrac{-100}{2*(-16)}\)

Mehek (mehek14):

2* - 16 = 32 -100/ 32 = -3.125

Mehek (mehek14):

now plug that in

OpenStudy (anonymous):

wait why is it negative? -/- is positive??

Mehek (mehek14):

oh sorry didn't see that :P

Mehek (mehek14):

-100/-32

Mehek (mehek14):

so 3.125

OpenStudy (anonymous):

oh okay so 3.125 do i plug it into the given function

Mehek (mehek14):

yes put it instead of t

OpenStudy (anonymous):

okay i got 159!

Mehek (mehek14):

so 159 is your answer :D

OpenStudy (anonymous):

thank you!! :)

Mehek (mehek14):

yw :)